OpenStudy (anonymous):

. A lamp is marked with a sale price of $23.80, which is 15% off of the regular price. What is the regular price? I know the answer, i want to know how to set up the equation to work the problem

OpenStudy (anonymous):

\[(1-.15)x=23.80\]

OpenStudy (anonymous):

where x is the regular price

OpenStudy (anonymous):

chmE, why did you take 1-15%

OpenStudy (anonymous):

because if you take 15% of the original price, you would then have to subtract it from the original price in order to get the reduced price. But if you just multiply it by 1-(percent off) you get the reduced price right away. I would take a calculator and prove this to yourself

OpenStudy (anonymous):

It's just like tax. If you need to add a 6% tax you could multiply the price by 6% then add it to the price to get the total with tax. But it is faster just to multiply by (1+.06) to get the final price right away

OpenStudy (anonymous):

i cant use a calculator on my test im going to take

OpenStudy (anonymous):

But you can now so that you can physically see that what I am saying is correct.

OpenStudy (anonymous):

so ur still paying 85%

OpenStudy (anonymous):

is that what the 85 is?

OpenStudy (anonymous):

ya. cuz if it is 15% off then you are really paying 85% of the original

OpenStudy (anonymous):

thats where the (1-.15) comes from

OpenStudy (anonymous):

mkk

OpenStudy (anonymous):

uhm this is probably really stupid but how do i divide with out a calc 23.80/.85 could u draw it

OpenStudy (anonymous):

you are going to have to do long division like in elementary days.I got rid of the decimals by multiplying the top and bottom by 100 first. This can also be seen like...\[\frac{23.80}{85/100} \rightarrow 23.80 \times \frac{100}{85} \rightarrow \frac{2380}{85}\]